The average length of time for students to register for summer classes at a certain college has been 50 minutes. A new registration procedure using modern computing machines is being tried. If a random sample of 35 students had an average registration time of 42 minutes with a standard deviation of 11.9 minutes under the new system, test the hypothesis that the population mean is now less than 50 minutes, using a level of signficance of 0.01. (critical value = +/- 2.33)

1
Expert's answer
2022-05-06T12:46:28-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge 50"

"H_a:\mu<50"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=34" degrees of fredom, and the critical value for a left-tailed test is "t_c = -2.44115."

The rejection region for this left-tailed test is "R = \{t: t < -2.44115\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{42-50}{11.9/\sqrt{35}}\approx-3.9772"

Since it is observed that "t = -3.9772< -2.44115=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=34" degrees of freedom, "t=-3.9772" is "p =0.000173," and since "p=0.000173<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 50, at the "\alpha = 0.01" significance level.